OpenStudy (anonymous):

I just dont understand these. 2 cos^2 θ − 5 cos θ + 2 = 0 sin^2 θ = 2 sin θ + 3 cos 2θ = 5 sin θ − 2 PLEASE HELP!!!

OpenStudy (anonymous):

Solve the given equation. (Enter your answers as a comma-separated list. Let k be any integer. Round terms to three decimal places where appropriate. If there is no solution, enter NO SOLUTION.)

OpenStudy (ranga):

I will show you the first, you can try the other two. 2 cos^2 θ − 5 cos θ + 2 = 0 This is basically a quadratic equation in cos(theta). Let t = cos(theta) 2t^2 - 5t + 2 = 0 2t^2 - 4t - t + 2 = 0 2t(t - 2) - 1(t - 2) = 0 (2t - 1)(t - 2) = 0 t = 1/2 or t = 2 Put t back as cos(theta) cos(theta) = 1/2 or cos(theta) = 2. The second can't be true because cos is always less than or equal to 1 and so you can discard that solution cos(theta) = 1/2 or theta = 60 degree or pi/3 radians. Since cos is cyclic, the function repeats itself every multiple of 2(pi). So theta = pi/3 +/- k(2(pi))

OpenStudy (anonymous):

ok thank you

OpenStudy (ranga):

you are welcome.

OpenStudy (anonymous):

Q= theta. ok so for the next one would it be sin^2Q-2sinQ+3?

OpenStudy (ranga):

sin^2 θ = 2 sin θ + 3 Let t = sin(theta) t^2 = 2t + 3 t^2 - 2t - 3 = 0. Solve for t. And then put back t as sin(theta) and then solve for theta. For problem 3) Use the double angle formula for cosines: cos(2A) = cos²A − sin²A = 2 cos²A − 1 = 1 − 2 sin²A The one that will be useful here is: cos(2A) = 1 − 2 sin²A Then follow the same procedure of temporarily replacing sin(theta) with t, solving for t and then putting back sin(theta) in the place of t. When you solve for t always be mindful that sine, cosine cannot be more than 1 or less than -1. So any solutions of t that is not in that range should be discarded. Also they ask you to use k as the integer. So any answer you get for theta you remember to add plus/minus k2(pi) where k = 0, 1, 2, 3, etc.

OpenStudy (anonymous):

Oh i was trying to do it step by step to make sure i was doing it right. Now i see how its -3